Neelam’s annual income is Rs. 288000. Her annual savings amount to Rs. 36000. The ratio of her savings to her expenditure is (A) 1 : 8 (B) 1 : 7 (C) 1 : 6 (D) 1 : 5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (B) is correct here.
From the question it is given that,
Neelam’s annual income is ₹288000
Her annual savings amount to₹36000
So, Neelam’s expenditure = 288000 –36000

=₹252000
Then, ratio of her savings to her expenditure = 36000/252000

= 36/252 Divide both numerator and denominator by 12.

= 3/21
Again, divide both numerator and denominator by 3.

= 1/7
Therefore, ratio of her savings to her expenditure = 1 : 7

Thanks